Q3. Frank is going to sell his 100,000 Google stock shares in three months. The current price is ...

Q3.

Frank is going to sell his 100,000 Google stock shares in three months. The current price is $5/share. To reduce the risk, he wants to buy a put option. Assume the price of Google share either goes up by 5% or down by 4% each month, though we can’t predict which event will happen exactly. The risk free interest rate is 1% each month. Gina sells Frank a put option. The strike price is 100,000 shares for $500,000.

a) What’s the price of this put option?

b) Describe Gina’s hedging portfolio two months after the option is issued.

Q4.

Carefully state the put-call parity. If put-call parity fails to hold, describe and confirm the arbitrage strategies.
Answer #1

Price   5       3month    1% per month
UP      5%           FSP = Future Spot Price  
Down 4%              
X or strike price   = 5              
Upward FSP   5.25              
Down FSP   4.8              
                  
Situation    FSP   Vc        
I                  5.25   -0.25          
II   4.8   0          
Change       0.45      -0.25          
                  
STRATEGY:                  
WRITE A CALL AND HOLD DELTA SHARES                  
                  
Calculation of Delta Shares                  
Delta shares = Change in VC        =   0.555555556 shares  
                       Change in FSP                  
Future value of portfolio/ maturity value of Portfolio                
Situation   FSP    Share Value   Vc          Portfolio value  
I               5.25      2.916666667   -0.25   2.666666667  
II                4.8      2.666666667   0           2.666666667  
Now we have created a risk less portfolio as we can see that the                  
in both situations the value of portfolio does not change !                  
                  
Present value of portfolio                 
The future value of portfolio is not influenced by FSP and is certain.                  
So to calculate Present Value, we can use Riskfree rate (ie 1%)                  
                  
So value (of portfolio) is discounted using continous discounting model                 
   2.66666667 / 1.01005      = 2.640133327      
                  
                  
Present value of delta shares                 
   Delta shares x Spot price              
   Delta shares       0.555555556      
   Spot Price                          5      
                       =      2.777777778     


Value of call                  
Value of Portfolio = Value of call +    Value of Delta shares              
2.640133 = Vc + 2.777778                  
Vc   0.137644451              
                  
Value of Put                
Put Call Parity equation                  
Vc + PVx = Vp + Vs                  
Present valueof strike price    5 / 1.01005       4.950249988      
                  
So Vp is                  
   0.087894439              

PUT CALL PARITY EQUATION

Vp+Vs = Vc+ PVx (where Vp is value of put, Vs is value of shares, Vc is value of call and PVx is Present value of spot price.

NOTE:

If the prices of the put and call options diverge so that this relationship does not hold good, an arbitrage opportunity exists, meaning that sophisticated traders can THEORETICALLY earn a risk-free profit. Such opportunities are uncommon and short-lived in liquid markets.